### v1

@@ 3.2 Extending to CAD and general QE
Extending to CAD’s projection phase. The framework Alg. 1 can be utilized in CAD’s projection phase with only minor modifications. We present the whole algorithm in Sect. 0.A.3 and here only explain the differences: (i) The two value functions $\mathcal{I}$ and $\mathcal{V}$ are defined as mappings from bags in $T$ to set of polynomials in $\mathbb{R}[x_{1},\dots,x_{n}]$, instead of atomic formulas; (ii) In the recursion step, for a forget bag $b$, we compute $\mathcal{V}(b)$ via $\mathcal{V}(b)\leftarrow\mathcal{I}(b)\cup\textsf{Proj}(\mathcal{V}(b.\texttt{children}),b.\texttt{forget})$, while for other types of bags, the computation of $\mathcal{V}(b)$ remains the same as in Alg. 1; and (iii) Finally, after applying the projection operation for all quantified variables, the CAD algorithm continues the projection phases for free variables, after which the lifting phase can start. The correctness of the framework follows from the fact that the algebraic operations in the projection operator (presented in Sect. 0.A.1) preserve locality, i.e., will not introduce new variables.
We shall note here that we cannot simply replace the invocation of the FME algorithm in Alg. 1 by the CAD algorithm, i.e., the projection phase and lifting phase of CAD must be separated. This is because the lifting phase produces a formula in disjunction normal form, on which we can not directly apply the dynamic programming framework.
@@
In fact, the variable ordering extracted via the above algorithm can be shown to be a perfect elimination ordering [^1] In graph theory, a perfect elimination ordering of a graph $G$ is an ordering of its vertices such that for every vertex $v$, the set consisting of $v$ and its neighbors that come after $v$ in the ordering forms a clique in $G$., say $(1,3,8,7,6,5,4,2)$ for our running example. Hence, it is possible that our variable ordering coincides with the one extracted from a chordal extension graph [li23jsc]. However, since we have exploited the sparsity information in building the tree decompositions, our variable orderings are usually superior for performing quantifier elimination on sparse problem instances. This will be further evidenced by our experimental results in Sect. 5. Moreover, our framework is stated for general quantifier elimination procedures and enables a fine-grained complexity analysis, explained in the following Sect. 4.
@@ 4 Complexity Analysis
Balanced tree-decomposition. The concept of $(\beta,\gamma)$-balanced tree decompositions [chatterjee18toplas] generalizes the notion of balanced tree decompositions that arise in the analysis of the space complexity of tree decomposition algorithms [bodlaender98joc, elberfeld10focs]. Informally, a $(\beta,\gamma)$-balanced tree-decomposition requires, for each bag $b$, the size of subtree $T_{b}$ decreases proportionally as $\textsf{Lv}(b)$ increases.
@@
The construction of a $(\beta,\gamma)$-balanced tree decomposition is elucidated in [chatterjee18toplas, Sec. 3], which gives the following theorem.
Theorem 4.1([chatterjee18toplas, Thm. 3.1]).
@@
In the following, we analyze the complexity of our frameworks for FME and CAD in Sect. 3 under the assumption that $k$ is a constant. For consistency [^2] Recall that the CAD’s projection phrase requires performing the projection operation for all variables, including the quantifier-free ones., we assume that $m=n$ in the input formula Eq. 1, i.e., there are no free variables.
@@
Complexity analysis for CAD. The CAD complexity is usually measured by the number of sign-invariant cells. Our analysis uses a similar approach as [li23jsc] but provides an improved bound based on the assumption of constant treewidth and the structure of balanced tree decompositions.
Before presenting the results, we introduce some tools to estimate the growth of the size of the polynomials in the projection phase. For a set of polynomials $P\subset\mathbb{R}[\mathbf{x}]$, we define the combined degree [mccallum85-thesis] of $P$ to be $\max_{1\leq i\leq n}\deg_{x_{i}}(\prod_{p\in P}p)$, where $\deg_{x_{i}}(\cdot)$ denotes the degree of a polynomial in a single variable $x_{i}$. A set of polynomials has the $(m,d)$-property if it can be partitioned into $m$ sets, such that each set has maximum combined degree $d$ [bradford16jsc, Def. 7]. It has been shown in [bradford16jsc, Lem. 11] that for a set $P\subset\mathbb{R}[\mathbf{x}]$ with the $(m,d)$-property, the set $\textsf{Proj}(P,x_{i})$ after projection has the $(M,2d^{2})$-property with $M=\lfloor\frac{(m+1)^{2}}{2}\rfloor$ for any $x_{i}\in\mathbf{x}$.
@@
The proof is inspired by the proof of [li23jsc, Thm. 11, Thm. 12]. Here we explain the main idea.
@@ 5 Implementation and Experiments
Computation Environment. We use the tool from [tamaki19jco-td] to compute tree decompositions, which are then processed by Python scripts. For LRA, we compare different heuristics based on a naive FME implementation in Python, while for NRA, we work on CAD implemented in Mathematica [mathematica]. All experiments were conducted on a Windows PC with i7-13700 CPU and 32GB RAM.
Existing Heuristics. For comparison, we briefly explain the most frequently used heuristic strategies for selecting the variable elimination ordering in FME and CAD. For FME, the baseline Random strategy makes $N=5$ trials according to different orderings that are sampled randomly and records the best result; the Greedy strategy chooses the next variable to be eliminated by minimizing the number of generated constraints. For CAD, the Brown’s heuristic, denoted Brown [brown04issac], chooses the next eliminated variable that has the lowest maximum degree across all polynomials in which it appears (and breaks ties with addition rules); the PEO strategy [li23jsc] extracts a perfect elimination ordering from the chordal extension of the primal graph; and we also compare with the default strategy in Maple [maple] to suggest variable ordering, denoted SVO.
@@
Our Implementation. In our practical implementation, we elected not to construct a balanced tree decomposition as described in Sect. 4. While the main advantage of the balanced tree decomposition lies in estimating the asymptotic complexity, it often performs suboptimally on problems of relatively small scale. Instead, we work with the tree decomposition computed by [tamaki19jco-td] and follow the approach detailed in Sect. 3.2 to extract the elimination order [^3] In our non-parallel setting, using the elimination order is functionally equivalent to running the dynamic programming framework, without the explicit computation of the value functions.. When traversing from top to bottom, we break ties using the aforementioned Greedy and Brown heuristics for FME and CAD, respectively.
@@
Benchmarks. As our algorithm targets dealing with LRA and NRA problems that are small enough to do quantifier elimination but also have low treewidth. This requirement rules out most existing standard benchmarks (an analysis on SMT-LIB benchmarks [barrett16smtlib] can be found in Sect. 0.A.4). Hence, for comparison, we generated a set of benchmarks with treewidth sparsity as follows: We construct benchmark inequality systems from graphs of prescribed treewidth $k$ by a two-stage procedure. We first generate a graph of treewidth $k$ by iteratively attaching nodes to a $(k+1)$-clique. Then, for each bag, we enumerate all monomials of total degree not exceeding the maximum degree ($1$ for LRA), select at least one highest-degree monomial, and randomly include other monomials with a probability ranging from 0.05 to 0.15, while ensuring that every scoped variable appears. For each set of hyperparameters, we randomly generated 10 examples for LRA and 1 example for NRA, because the randomly selected monomials significantly impact the CAD efficiency, and the average runtime is usually dominated by the worst instance.
@@
Results. Table 1 and Table 2 present the results of FME and CAD, respectively. On LRA benchmarks, our strategy consistently yielded the lowest average inequality count and average runtime, achieving the best result in 4 out of 6 sets. The Greedy strategy showed a very similar performance with the runtime difference being negligible, but was significantly slower on ID 3 and 5. The Random strategy performed poorly, generating an extremely high number of inequalities, often exceeding $10^{7}$ for most instances. On NRA benchmarks, our strategy also demonstrates superior performance in all but one benchmark, in terms of both the number of cells generated and the runtime. For benchmark ID 9 and ID 12, it significantly outperforms all other methods. The only instance where our strategy did not have the lowest cell count was ID 10, where PEO was superior. The results answer our research question affirmatively, indicating that our proposed strategy is effective and more efficient in dealing with problems with low treewidth, compared to the existing heuristics. More experiments on benchmarks from [li23jsc] are available at Sect. 0.A.5.
@@ 6 Conclusion
Data Availability Statement. The code for our experiments is available at https://github.com/Jiyu-ISCAS/qe-treewidth-artifact.
@@ 0.A.1 CAD Algorithm
In this part, we provide a formal description of the CAD algorithm. The notations and definitions mostly follow Jirstrand’s introductory paper [jirstrand95]. Given a set of polynomials in $\mathcal{F}\subset\mathbb{R}[x_{1},\dots,x_{n}]$, the (CAD) algorithm decomposes the $\mathbb{R}^{n}$ space into finitely many sign-invariant regions, called cells. After decomposition, one can determine the sign of the polynomials in each cell using sampled points in this cell.
@@
Definition 8(McCallum’s Projection Operator [mccallum88jsc]).
@@
In each step, the zero sets of the polynomials in $\mathcal{F}_{i-1}$ is the projection of “significant points” of the zero set of polynomial in $\mathcal{F}_{i}$, i.e. self crossings, isolated points, vertical tangent points etc.
@@ 0.A.3 Extensions of the Dynamic Programming Framework
3 Construct the primal graph $G_{\Phi}=(V_{\Phi},E_{\Phi})$ of $\Phi$;
4 Compute a nice tree decomposition $T=(\mathcal{B},E_{T})$ of $G_{\Phi}$;
5 $\triangleright$ Initialize $\mathcal{I}$ from top down.
6 $P\subset\mathbb{R}[\mathbf{x}]\leftarrow$ the set of polynomials in $\Phi$;
7 Initialize an empty queue of bags $q=\langle\rangle$;
8 $q.\texttt{enqueue}(r_{T})$ ;
9 while $q$ is not empty do
10 $b\leftarrow q.\texttt{dequeue}()$;
11 $q.\texttt{enqueue}(b.\texttt{children})$;
12 if $b.\textnormal{{type}}=\textsf{leaf}$ then
13    $\mathcal{I}(b)\leftarrow\emptyset$;
14
15 else
16    $\mathcal{I}(b)\leftarrow\{p\in P\mid\texttt{var}(p)\subseteq b\}$;
17    $P\leftarrow P\setminus\mathcal{I}(b)$;
18
19   end if
20
21 end while
$2\triangleright$ Recursively compute $\mathcal{V}$ from bottom up.
23 $q.\texttt{enqueue}(\mathcal{B}.\texttt{leaves})$ ;
24 while $q$ is not empty do
25 $b\leftarrow q.\texttt{dequeue}()$;
26 $q.\texttt{enqueue}(b.\texttt{parent})$;
27 switch $b.\textnormal{{type}}$ do
28    case join do $\mathcal{V}(b)\leftarrow\mathcal{I}(b)\cup\mathcal{V}(b.\texttt{leftchild})\cup\mathcal{V}(b.\texttt{rightchild})$;
29    case introduce do $\mathcal{V}(b)\leftarrow\mathcal{I}(b)\cup\mathcal{V}(b.\texttt{children})$;
30    case forget do $\mathcal{V}(b)\leftarrow\mathcal{I}(b)\cup\textsf{Proj}(\bigwedge\mathcal{V}(b.\texttt{children}),b.\texttt{forget})$;
31    case leaf do $\mathcal{V}(b)\leftarrow\mathcal{I}(b)$;
32
33   end switch
34
35 end while
36return $\bigcup\mathcal{V}(r_{T})$;
[…]
@@ 17    $P\leftarrow P\setminus\mathcal{I}(b)$;
1
21exConstruct the primal graph $G_{\Phi}=(V_{\Phi},E_{\Phi})$ of $\Phi$;
3 Compute a (nice) tree decomposition $T=(\mathcal{B},E_{T})$ of $G_{\Phi}$;
4 $\triangleright$ Extract a variable elimination order from $T$.
5 Initialize an empty queue of variables $order=\langle\rangle$;
6 Initialize an empty queue of bags $q=\langle\rangle$;
7 $q.\texttt{enqueue}(r_{T})$;
8 while $q$ is not empty do
9 $b\leftarrow\texttt{dequeue}(q)$;
10 $\texttt{enqueue}(q,b.\texttt{children})$;
11 if $b$ is root then
12    $order.\texttt{enqueue}(b)$;
13
14 else
15    $order.\texttt{enqueue}(b\setminus b.\textsf{parent})$
16   end if
17
18 end while
[…]
20 // reverse the ordering
@@ 0.A.4 Sparsity in SMT-LIB benchmarks
We examined the QF_LRA and QF_NRA categories from the SMT-LIB benchmark library [barrett16smtlib]. Using pysmt, we successfully parsed 1753 LRA instances and 9249 NRA instances. Among the parsed instances, only 337 LRA instances and 807 NRA instances have non-complete primal graphs. We first analyzed the treewidth of these instances, the results are presented in Table 3 and Table 4. However, since many instances contain hundreds or more variables, it is impractical to perform a full tree decomposition for every instance (e.g., within a 10-minute timeout). Therefore, we also examined the edge density sparsity, defined by $|E|/\binom{|V|}{2}$, where $E$ and $V$ denote the number of edges and vertices of a primal graph, respectively. The results are presented in Table 5 and Table 6. From these results, we found that SMT-LIB benchmarks are generally too large and too dense for demonstrating our algorithms.
Table 3: Treewidth distribution among 337 sparse instances in the QF_LRA dateset
@@
Table 4: Treewidth distribution among 807 sparse instances in the QF_NRA dateset
@@
Table 5: Edge density among 337 sparse instances in the QF_LRA dateset
@@
Table 6: Edge density among 807 sparse instances in the QF_NRA dateset

### v2

@@ 3.2 Extending to CAD and General QE
Extending to CAD’s projection phase. The framework Alg. 1 can be utilized in CAD’s projection phase with only minor modifications. We present the whole algorithm in the full version Sect. 0.A.3 and only explain the differences here: (i) The two value functions $\mathcal{I}$ and $\mathcal{V}$ are defined as mappings from bags in $T$ to set of polynomials in $\mathbb{R}[x_{1},\dots,x_{n}]$, instead of atomic formulas; (ii) In the recursion step, for a forget bag $b$, we compute $\mathcal{V}(b)$ via $\mathcal{V}(b)\leftarrow\mathcal{I}(b)\cup\textsf{Proj}(\mathcal{V}(b.\texttt{children}),b.\texttt{forget})$, while for other types of bags, the computation of $\mathcal{V}(b)$ remains the same as in Alg. 1; and (iii) Finally, after applying the projection operation for all quantified variables, the CAD algorithm continues the projection phases for free variables, after which the lifting phase can start. The correctness of the framework follows from the fact that the algebraic operations in the projection operator (presented in Sect. 0.A.1) preserve locality, i.e., will not introduce new variables. We shall note here that we cannot simply replace the invocation of the FME algorithm in Alg. 1 by the CAD algorithm, i.e., the projection phase and lifting phase of CAD must be separated. This is because the lifting phase produces a formula in disjunction normal form, on which we can not directly apply the dynamic programming framework.
@@
In fact, the variable ordering extracted via the above algorithm can be shown to be a perfect elimination ordering [^1] In graph theory, a perfect elimination ordering of a graph $G$ is an ordering of its vertices such that for every vertex $v$, the set consisting of $v$ and its neighbors that come after $v$ in the ordering forms a clique in $G$., say $(1,3,8,7,6,5,4,2)$ for our running example. Hence, it is possible that our variable ordering coincides with the one extracted from a chordal extension graph [54]. However, since we have exploited the sparsity information in building the tree decompositions, our variable orderings are usually superior for performing quantifier elimination on sparse problem instances. This will be further evidenced by our experimental results in Sect. 5. Moreover, our framework is stated for general quantifier elimination procedures and enables a fine-grained complexity analysis, explained in the following Sect. 4.
@@ 4 Complexity Analysis
Balanced tree-decomposition. The concept of $(\beta,\gamma)$-balanced tree decompositions [17] generalizes the notion of balanced tree decompositions that arise in the analysis of the space complexity of tree decomposition algorithms [6, 30]. Informally, a $(\beta,\gamma)$-balanced tree-decomposition requires, for each bag $b$, the size of subtree $T_{b}$ decreases proportionally as $\textsf{Lv}(b)$ increases.
@@
The construction of a $(\beta,\gamma)$-balanced tree decomposition is elucidated in [17, Sec. 3], which gives the following theorem.
Theorem 4.1([17, Thm. 3.1]).
@@
In the following, we analyze the complexity of our frameworks for FME and CAD in Sect. 3 under the assumption that $k$ is a constant. For consistency [^2] Recall that the CAD’s projection phase requires performing the projection operation for all variables, including the quantifier-free ones., we assume that $m=n$ in the input formula Eq. 1, i.e., there are no free variables.
@@
Complexity analysis for CAD. The CAD complexity is usually measured by the number of sign-invariant cells. Our analysis uses a similar approach as [54] but provides an improved bound based on the assumption of constant treewidth and the structure of balanced tree decompositions.
Before presenting the results, we introduce some tools to estimate the growth of the size of the polynomials in the projection phase. For a set of polynomials $P\subset\mathbb{R}[\mathbf{x}]$, we define the combined degree [59] of $P$ to be $\max_{1\leq i\leq n}\deg_{x_{i}}(\prod_{p\in P}p)$, where $\deg_{x_{i}}(\cdot)$ denotes the degree of a polynomial in a single variable $x_{i}$. A set of polynomials has the $(m,d)$-property if it can be partitioned into $m$ sets, such that each set has maximum combined degree $d$ [10, Def. 7]. It has been shown in [10, Lem. 11] that for a set $P\subset\mathbb{R}[\mathbf{x}]$ with the $(m,d)$-property, the set $\textsf{Proj}(P,x_{i})$ after projection has the $(M,2d^{2})$-property with $M=\lfloor\frac{(m+1)^{2}}{2}\rfloor$ for any $x_{i}\in\mathbf{x}$.
@@
The proof is inspired by the proof of [54, Thm. 11, Thm. 12]. Here we explain the main idea.
@@ 5 Implementation and Experiments
Computation Environment. We use the tool from [72] to compute tree decompositions, which are then processed by Python scripts. For LRA, we compare different heuristics based on a naive FME implementation in Python, while for NRA, we work on CAD implemented in Mathematica [45]. All experiments were conducted on a Windows PC with i7-13700 CPU and 32GB RAM.
Existing Heuristics. For comparison, we briefly explain the most frequently used heuristic strategies for selecting the variable elimination ordering in FME and CAD. For FME, the baseline Random strategy makes $N=5$ trials according to different orderings that are sampled randomly and records the best result; the Greedy strategy chooses the next variable to be eliminated by minimizing the number of generated constraints. For CAD, the Brown’s heuristic, denoted Brown [14], chooses the next eliminated variable that has the lowest maximum degree across all polynomials in which it appears (and breaks ties with addition rules); the PEO strategy [54] extracts a perfect elimination ordering from the chordal extension of the primal graph; and we also compare with the default strategy in Maple [57] to suggest variable ordering, denoted SVO.
@@
Our Implementation. In our practical implementation, we elected not to construct a balanced tree decomposition as described in Sect. 4. While the main advantage of the balanced tree decomposition lies in estimating the asymptotic complexity, it often performs suboptimally on problems of relatively small scale. Instead, we work with the tree decomposition computed by [72] and follow the approach detailed in Sect. 3.2 to extract the elimination order [^3] In our non-parallel setting, using the elimination order is functionally equivalent to running the dynamic programming framework, without the explicit computation of the value functions.. When traversing from top to bottom, we break ties using the aforementioned Greedy and Brown heuristics for FME and CAD, respectively.
@@
Benchmarks. As our algorithm targets dealing with LRA and NRA problems that are small enough to do quantifier elimination but also have low treewidth. This requirement rules out most existing standard benchmarks (an analysis on SMT-LIB benchmarks [5] can be found in the full version Sect. 0.A.4. Hence, for comparison, we generated a set of benchmarks with treewidth sparsity as follows: We construct benchmark inequality systems from graphs of prescribed treewidth $k$ by a two-stage procedure. We first generate a graph of treewidth $k$ by iteratively attaching nodes to a $(k+1)$-clique. Then, for each bag, we enumerate all monomials of total degree not exceeding the maximum degree ($1$ for LRA), select at least one highest-degree monomial, and randomly include other monomials with a probability ranging from 0.05 to 0.15, while ensuring that every scoped variable appears. For each set of hyperparameters, we randomly generated 10 examples for LRA and 1 example for NRA, because the randomly selected monomials significantly impact the CAD efficiency, and the average runtime is usually dominated by the worst instance.
@@
Results. Table 1 and Table 2 present the results of FME and CAD, respectively. On LRA benchmarks, our strategy consistently yielded the lowest average inequality count and average runtime, achieving the best result in 4 out of 6 sets. The Greedy strategy showed a very similar performance with the runtime difference being negligible, but was significantly slower on ID 3 and 5. The Random strategy performed poorly, generating an extremely high number of inequalities, often exceeding $10^{7}$ for most instances. On NRA benchmarks, our strategy also demonstrates superior performance in all but one benchmark, in terms of both the number of cells generated and the runtime. For benchmark ID 9 and ID 12, it significantly outperforms all other methods. The only instance where our strategy did not have the lowest cell count was ID 10, where PEO was superior. The results answer our research question affirmatively, indicating that our proposed strategy is effective and more efficient in dealing with problems with low treewidth, compared to the existing heuristics. More experiments on benchmarks from [54] are available at Sect. 0.A.5.
@@ 6 Conclusion
Data Availability Statement. The code for our experiments is available at https://doi.org/10.5281/zenodo.18150640.
@@ 0.A.1 CAD Algorithm
In this part, we provide a formal description of the CAD algorithm. The notations and definitions mostly follow Jirstrand’s introductory paper [48]. Given a set of polynomials in $\mathcal{F}\subset\mathbb{R}[x_{1},\dots,x_{n}]$, the (CAD) algorithm decomposes the $\mathbb{R}^{n}$ space into finitely many sign-invariant regions, called cells. After decomposition, one can determine the sign of the polynomials in each cell using sampled points in this cell.
@@
Definition 8(McCallum’s Projection Operator [60]).
@@
In each step, the zero sets of the polynomials in $\mathcal{F}_{i-1}$ are the projection of “significant points” of the zero set of polynomial in $\mathcal{F}_{i}$, i.e. self crossings, isolated points, vertical tangent points etc.
@@ 0.A.3 Extensions of the Dynamic Programming Framework
Construct the primal graph $G_{\Phi}=(V_{\Phi},E_{\Phi})$ of $\Phi$;
Compute a nice tree decomposition $T=(\mathcal{B},E_{T})$ of $G_{\Phi}$;
$\triangleright$ Initialize $\mathcal{I}$ from top down.
$P\subset\mathbb{R}[\mathbf{x}]\leftarrow$ the set of polynomials in $\Phi$;
Initialize an empty queue of bags $q=\langle\rangle$;
$q.\texttt{enqueue}(r_{T})$ ;
while $q$ is not empty do
 $b\leftarrow q.\texttt{dequeue}()$;
 $q.\texttt{enqueue}(b.\texttt{children})$;
 if $b.\textnormal{{type}}=\textsf{leaf}$ then
    $\mathcal{I}(b)\leftarrow\emptyset$;
 else
    $\mathcal{I}(b)\leftarrow\{p\in P\mid\texttt{var}(p)\subseteq b\}$;
    $P\leftarrow P\setminus\mathcal{I}(b)$;
   end if
end while
$\triangleright$ Recursively compute $\mathcal{V}$ from bottom up.
$q.\texttt{enqueue}(\mathcal{B}.\texttt{leaves})$ ;
while $q$ is not empty do
 $b\leftarrow q.\texttt{dequeue}()$;
 $q.\texttt{enqueue}(b.\texttt{parent})$;
 switch $b.\textnormal{{type}}$ do
    case join do $\mathcal{V}(b)\leftarrow\mathcal{I}(b)\cup\mathcal{V}(b.\texttt{leftchild})\cup\mathcal{V}(b.\texttt{rightchild})$;
    case introduce do $\mathcal{V}(b)\leftarrow\mathcal{I}(b)\cup\mathcal{V}(b.\texttt{children})$;
    case forget do $\mathcal{V}(b)\leftarrow\mathcal{I}(b)\cup\textsf{Proj}(\bigwedge\mathcal{V}(b.\texttt{children}),b.\texttt{forget})$;
    case leaf do $\mathcal{V}(b)\leftarrow\mathcal{I}(b)$;
   end switch
end while
return $\bigcup\mathcal{V}(r_{T})$;
$\triangleright$ After this algorithm finishes, the CAD algorithm continuous the projection phase for free variables, until only one variable is left. Then the lifting phase begins.
@@
1exConstruct the primal graph $G_{\Phi}=(V_{\Phi},E_{\Phi})$ of $\Phi$;
Compute a (nice) tree decomposition $T=(\mathcal{B},E_{T})$ of $G_{\Phi}$;
$\triangleright$ Extract a variable elimination order from $T$.
Initialize an empty queue of variables $order=\langle\rangle$;
Initialize an empty queue of bags $q=\langle\rangle$;
$q.\texttt{enqueue}(r_{T})$;
while $q$ is not empty do
 $b\leftarrow\texttt{dequeue}(q)$;
 $\texttt{enqueue}(q,b.\texttt{children})$;
 if $b$ is root then
    $order.\texttt{enqueue}(b)$;
 else
    $order.\texttt{enqueue}(b\setminus b.\textsf{parent})$
   end if
end while
$order\leftarrow order.\texttt{reverse}$ ;
// reverse the ordering
@@ 0.A.4 Sparsity in SMT-LIB Benchmarks
We examined the QF_LRA and QF_NRA categories from the SMT-LIB benchmark library [5]. Using pysmt, we successfully parsed 1753 LRA instances and 9249 NRA instances. Among the parsed instances, only 337 LRA instances and 807 NRA instances have non-complete primal graphs. We first analyzed the treewidth of these instances, the results are presented in Table 3 and Table 4. However, since many instances contain hundreds or more variables, it is impractical to perform a full tree decomposition for every instance (e.g., within a 10-minute timeout). Therefore, we also examined the edge density sparsity, defined by $|E|/\binom{|V|}{2}$, where $E$ and $V$ denote the number of edges and vertices of a primal graph, respectively. The results are presented in Table 5 and Table 6. From these results, we found that SMT-LIB benchmarks are generally too large and too dense for demonstrating our algorithms.
Table 3: Treewidth distribution among 337 sparse instances in the QF_LRA dataset
@@
Table 4: Treewidth distribution among 807 sparse instances in the QF_NRA dataset
@@
Table 5: Edge density among 337 sparse instances in the QF_LRA dataset
@@
Table 6: Edge density among 807 sparse instances in the QF_NRA dataset